### v1

# Generative or Discriminative? Revisiting Text Classification in the Era of Transformers

Abstract
The comparison between discriminative and generative classifiers has intrigued researchers since Efron (1975)’s seminal analysis of logistic regression versus discriminant analysis. While early theoretical work established that generative classifiers exhibit lower sample complexity but higher asymptotic error in simple linear settings, these trade-offs remain unexplored in the transformer era. We present the first comprehensive evaluation of modern generative and discriminative architectures—Auto-regressive, Masked Language Modeling, Discrete Diffusion, and Encoders for text classification. Our study reveals that the classical “two regimes" phenomenon manifests distinctly across different architectures and training paradigms. Beyond accuracy, we analyze sample efficiency, calibration, noise robustness, and ordinality across diverse scenarios. Our findings offer practical guidance for selecting the most suitable modeling approach based on real-world constraints such as latency and data limitations.
Generative or Discriminative?
Revisiting Text Classification in the Era of Transformers
Siva Rajesh Kasa, Karan Gupta, Sumegh Roychowdhury, Ashutosh Kumar, Yaswanth Biruduraju,
Santhosh Kumar Kasa, Nikhil Priyatam Pattisapu, Arindam Bhattacharya,
Shailendra Agarwal, Vijay huddar Amazon Inc. kasasiva,karaniis,sumegr@amazon.com
1 Introduction
Text Classification (TC), a fundamental task in Natural Language Processing (NLP), encompasses various applications such as Sentiment Analysis, Topic Classification, and Emotion Detection. Since the emergence of transformer architectures, the field has been dominated by discriminative classifiers that leverage token embeddings (e.g., the [CLS] token in BERT Devlin et al. (2019)). These models directly learn the conditional probability distribution $P_{\theta}(y|X)$, where $X$ denotes the input text and $y$ represents the ground truth label. However, as these discriminative models grow larger, they require increasingly large amounts of labeled data to achieve optimal performance, making them impractical in many real-world scenarios where labeled data is scarce or expensive to obtain (Zheng et al., 2023). On the other hand, generative classifiers, which model the joint distribution $P_{\theta}(X,y)$, are known to work better in low-data settings, giving rise to the classical ‘two-regimes’ phenomenon for classification (Ng and Jordan, 2001; Yogatama et al., 2017; Zheng et al., 2023). This advantage stems from their ability to learn underlying data distributions rather than just decision boundaries, allowing them to make better use of limited training examples. The inherent data efficiency of generative approaches, combined with recent advances in generative modeling such as Discrete Diffusion (Lou et al., 2024), motivates us to revisit the classical discriminative versus generative debate in the context of TC with Transformer-based architectures.
Prior research on generative classifiers has largely focused on non-textual tabular data, utilizing linear models such as Linear Discriminant Analysis (Efron, 1975) and Naive Bayes (Ng and Jordan, 2001). While Yogatama et al. (2017) extended this analysis to neural architectures using RNNs and LSTMs (Hochreiter and Schmidhuber, 1997) for the TC task and found similar conclusions about generative advantages in low-data regimes, their study predated the transformer era. Modern NLP has seen the emergence of various successful transformer-based generative modeling paradigms, including auto-regressive (AR) models like GPT (Radford et al., 2018) that maximize likelihood directly, Discrete Diffusion models (Lou et al., 2024) that learn through iterative denoising, and masked language models (MLM) (Devlin et al., 2019) that optimize pseudo-likelihood Wang and Cho (2019). These approaches offer different trade-offs in terms of computational efficiency, sample complexity, and modeling flexibility. However, a systematic comparison of these paradigms for text classification remains unexplored, particularly in the context of varying model sizes and real-world deployment considerations.Our work fills this gap by providing a practitioner-oriented study that evaluates these approaches not just on classification accuracy, but also on crucial deployment metrics including different model scales, robustness to input perturbations, reliability of output probabilities through calibration analysis, and preservation of ordinal relationships between classes. This comprehensive evaluation aims to provide concrete guidance for choosing between different generative and discriminative approaches based on specific deployment constraints and requirements. We strategically focus on widely available public benchmark datasets for reproducibility purposes. Following Li et al. (2025) and Yogatama et al. (2017), our study evaluates all models trained from scratch, rather than relying on pre-trained weights, providing crucial insights for practitioners working with domain-specific data (Huang et al., 2019) or in resource-constrained environments (Martin et al., 2022). This approach helps isolate the confounding effects of the pre-training corpus (Razeghi et al., 2022) from other factors such as the modeling approach and size, which we evaluate. The combined effects can be explored in future work.
Our main contributions include the following: (a) We present the first large-scale comparative study of two major classification approaches - Discriminative (Encoder) and Generative (Text Diffusion, AR, MLM) on 9 popular classification benchmark datasets, which is a first of its kind in the transformer era. Our study reveals a more nuanced interplay between model size and sample complexity than the previously known “two regimes” phenomenon. (b) We provide comprehensive analyses across multiple dimensions including model scaling behavior, sample efficiency, and performance in low-resource settings with models trained from scratch. We also introduce novel evaluation perspectives by examining ordinal relationships between classes, output calibration and robustness to input noise, offering insights beyond traditional classification metrics. (c) Finally, we provide practical recommendations in Section 6 on selecting the appropriate model for deployment in various real-world scenarios.
Figure 1: [Best viewed in color] Illustration of different modeling paradigms (ENC: Encoder-based classification, MLM: Masked Language Modeling, AR: Auto-Regressive Model, DIFF: Discrete Text Diffusion).
2 Related Work
Generative and Discriminative Models for Classification. The comparison between generative and discriminative classifiers originated with Efron (1975)’s analysis of logistic regression and discriminant analysis. Building on this foundation, Ng and Jordan (2001) examined naive Bayes and logistic regression, establishing the fundamental trade-off between generative models’ faster learning rate and discriminative models’ lower asymptotic error. Their theoretical analysis heavily depends on linearity and independence assumptions. However, subsequent work by Xue and Titterington (2008) challenged these findings through empirical studies and asymptotic analysis of statistical efficiency. Yogatama et al. (2017) provided the first empirical study of discriminative vs generative models for TC with neural architectures using LSTMs. They maximize the joint probability $P(X,y)=P(X|y)P(y)$ by concatenating the label $y$ text at the beginning of the input text $X$ and maximizing the class conditional likelihood i.e. $P(X\mid y)=\prod_{t=1}^{T}p\left(x_{t}\mid\boldsymbol{x}_{<t},y\right)$. The final predicted label is obtained by $\hat{y}=\operatorname*{\texttt{argmax}}_{y}P(X|y)P(y)$. They found that generative LSTMs have better accuracy than their discriminative counterparts at low-sample regimes. Further, they noted that the neural generative LSTMs are generally better than baseline generative models with stronger independence assumptions (e.g. naive Bayes, Kneser–Ney Bayes (Ney et al., 1994; Teh, 2006)). Next, the work by Zheng et al. (2023) has extended the theoretical understanding of generative classifiers to multi-class and non-linear settings. More recent studies (Li et al., 2025; Stanley et al., 2025) have found that generative classifiers tend to avoid shortcut learning and exhibit greater robustness to distribution shifts.
While prior studies provide valuable insights, the landscape of NLP has evolved dramatically with the advent of novel transformer-based generative paradigms such as Auto-Regressive (AR) models (Radford et al., 2018) and Discrete Diffusion models (Lou et al., 2024). Our work extends beyond these previous comparisons by conducting the first comprehensive evaluation of modern transformer-based generative and discriminative classifiers for TC. While previous works primarily focused on classification accuracy and sample complexity, we examine multiple dimensions that are crucial for real-world deployments. For instance, Yogatama et al. (2017) initial work with neural architectures was limited to a fixed model size, leaving open questions about how the generative-discriminative trade-off varies with model capacity and computational budget—questions that have become increasingly relevant in the era of large language models. Similarly, though Zheng et al. (2023) provided theoretical insights for multi-class settings, their analysis did not address practical considerations like calibration quality or preservation of ordinal relationships between classes.
Pseudo-Generative Models. Recent work Sahoo et al. (2024) highlights a natural connection between Discrete Text Diffusion Lou et al. (2024) and the Masked Language Modeling (MLM) objective in BERT Devlin et al. (2019), showing that the diffusion objective can be expressed as a weighted sum of MLM losses. Using transformer encoder models, this approach achieves likelihood bounds comparable to or better than those in Lou et al. (2024). Motivated by this, we include vanilla MLM as a baseline for text classification. While MLM has typically served as a pretraining objective followed by fine-tuning Liu et al. (2019), there has been little systematic study of its direct use for classification. Although MLM does not explicitly model $P(X|y)$, it estimates $P(x_{m}|x_{\setminus m})$, where $x_{m}$ is a masked token and $x_{\setminus m}$ represents all other tokens. This approximates the pseudo-likelihood of $P(X,y)$ when modeled over the corpus (Wang and Cho, 2019). We therefore classify MLM as a pseudo-generative model.
Also, traditional generative classifiers aim to model $P(X|y)$ by prepending the label token. However, recent work Li et al. (2025) shows that appending the label at the end—though not strictly modeling $P(X|y)$—can yield better in-distribution performance. This setup also enables efficient inference, requiring only a single forward pass to predict the label, unlike traditional generative models that need #$\#_{\text{label}}$ forward passes. These benefits motivate the inclusion of such pseudo-generative models in our benchmarks. Notably, these approaches involve minimal changes to standard transformer architectures—typically just altering label placement or the loss function—while preserving the core model design. This allows for fair comparisons using widely available implementations accessible to practitioners.
We also acknowledge a separate class of hybrid generative-discriminative models, where some subset of parameters are trained generatively and others discriminatively Raina et al. (2003); McCallum et al. (2006); Hayashi (2025). However, we exclude them from our study, as their architectural differences hinder fair comparison with fully generative or discriminative models, placing them outside the scope of this work.
Relation to Multi-task Learning. Learning $\log P(X,y)$ jointly, when factored as $\log P(X)+\log P(y|X)$ (or $\log P(y)+\log P(X|y)$) can be viewed as a multi-task learning setup, where unsupervised learning of $\log P(X)$ ($\log P(Y)$) and supervised learning of $\log P(Y|X)$ ($\log P(X|Y)$) represent two different but related tasks. This connection is supported by empirical results showing that unsupervised pre-training helps downstream supervised tasks (Erhan et al., 2010). As demonstrated by Wu et al. (2020); Hu et al. (2023), when model capacity is sufficiently large, such multi-task learning setups tend to be more successful - the model has enough capacity to perform well on both the unsupervised and supervised objectives. However, with limited model capacity, there are inherent trade-offs between the tasks, leading to challenges in jointly optimizing for both $P(X)$ and $P(y|X)$ (or $P(y)$ and $P(X|y)$). This insight motivates us to conduct a systematic study examining the relationship between model capacity and the performance of discriminative vs generative classifiers - an analysis that has not been previously undertaken in the literature.
Refer to Appendix A for further related works review on Discrete Diffusion, Robustness to Noise, Ordinality & Calibration.
3 Methodology
We approach the problem of label classification by leveraging two popular language modeling paradigms: (a) Generative - Discrete Diffusion models, Auto-regressive models (AR), and Masked Language Models (MLM) & (b) Discriminative - Encoder-based transformer models. Note that, for brevity, we use the term “generative” from this point onward to also include the pseudo-generative baselines. Let $\mathcal{D}=\{(X_{i},y_{i})\}_{i=1}^{N}$ denote the dataset where $X_{i}$ is the input text and $y_{i}\in\mathcal{Y}$ is the corresponding label from a finite set of classes $\mathcal{Y}$. Generative models tend to learn the joint data distribution $P(X,y)$ first and then try to infer the label using the marginals, whereas Discriminative models directly learn the conditional distribution $P(y|X)$. Note that each $X_{i}=x_{i}^{1}\ldots x_{i}^{n}$, where $x_{i}^{j}$ is a token from the associated vocabulary $\mathcal{V}$.
3.1 Discriminative Model for Classification
(1) Encoder-based classification (ENC): A Transformer encoder Vaswani et al. (2017) $f_{\theta}$ encodes the input as $h_{i}=f_{\theta}(X_{i})$ as a $d$-dimensional embedding, followed by a linear classifier head $W\in\mathbb{R}^{|\mathcal{Y}|\times d}$ which is the standard discriminative learning setup:
$\hat{y}_{i}=\texttt{softmax}(Wh_{i}),\quad\mathcal{L}_{\text{enc}}=-\sum_{i=1}^{N}\log P(y_{i}|X_{i})$ (1)
where $\mathcal{L}_{\text{enc}}$ is the cross-entropy based objective for training the encoder model.
3.2 Generative Models for Classification
(2) Masked Language Modeling (MLM): During training, we mask 15% of tokens in input sequence $X_{i}=x_{i}^{1}\ldots x_{i}^{n}$ following Devlin et al. (2019) and predict them using unmasked bi-directional context. Wang and Cho (2019) show that the MLM objective stochastically captures the pseudo-loglikelihood which makes it similar to a denoising autoencoder Vincent et al. (2010). Hence, we consider MLM under the generative family of models. Formally, the objective is:
$\mathcal{L}_{\text{mlm}}=-\sum_{i=1}^{N}\sum_{j\in\mathcal{M}_{i}}\log P(x_{i}^{j}|X_{i}^{\setminus j})$ (2)
where $\mathcal{M}_{i}$ is the set of masked positions and $X_{i}^{\setminus j}$ denotes the unmasked input with only token at position $j$ masked. At inference, we use the template:
[…]
and predict the masked label token. The output vocabulary is restricted to the label token set $\mathcal{V}_{\mathcal{Y}}$.
(3) Auto-regressive modeling (AR): Following Radford et al. (2018), we train a causal generative model to minimize the next-token prediction loss over the entire label + input sequence:
$\mathcal{L}_{\text{gpt}}=-\sum_{i=1}^{N}\sum_{j=1}^{L_{i}}\log P(x_{i}^{j}|y,x_{i}^{1},\ldots,x_{i}^{j-1})$ (3)
where $L_{i}$ is the length of the $i$-th sequence. At inference time, we perform one forward pass per candidate label $y\in\mathcal{V}_{\mathcal{Y}}$ by prepending it to the input $X$, and compute the log-likelihood. The predicted label is then obtained as $\arg\max_{y\in\mathcal{V}_{\mathcal{Y}}}\log P(X\mid y)$. In ARpseudo (refer pseudo-generative models in Section 2) the label is appended at the end instead of the beginning and only one forward pass is required to generate the predicted label token $y$. Note that label placement is only relevant for causal generative architectures (like AR) with a left-to-right attention structure. For bidirectional (pseudo-)generative models like MLM or DIFF, it has no theoretical impact.
(4) Text Diffusion (DIFF): For each input-label pair $(X_{i},y_{i})$, we first create a template:
$X_{i}=X_{i}\;\texttt{[SEP]}\;\texttt{"The label is"}\;y_{i}\;\texttt{.}$
where each template is a sequence $X_{i}={x}_{i}^{1}\ldots{x}_{i}^{L_{i}}$ with tokens ${x}_{i}^{j}\in\mathcal{V}$.
Similar to how diffusion models gradually add noise to images, our forward process gradually corrupts text by converting tokens to pure noise (here [MASK]). Following Lou et al. (2024), we define the forward process through discrete transition matrices $Q_{t}$ following a continuous markov process (see eq. 4). This process occurs at different timesteps $t\in[0,T]$, where each token position is independently corrupted, starting from the original text and progressively moving towards a completely masked sequence.
$\frac{dp_{t}}{dt}=Q_{t}\,p_{t},\quad\text{with}\quad p_{0}=p_{\text{data}}$ (4)
The reverse process learns to reconstruct the original text by predicting what token should replace each [MASK] symbol. This is done by learning score ratios $s_{\theta}(x,t)_{z}=\frac{p_{t}(z)}{p_{t}(x)}$ where $x,z$ are tokens from $\mathcal{V}$ and modeling the reverse process Sun et al. (2022) as:
$\frac{dp_{T-t}}{dt}=s_{\theta}(x,t)_{z}Q_{T-t}\,p_{T-t}$ (5)
Denoising Score Entropy (DSE) is used for training the score model in a manner that ensures several desired properties for $s_{\theta}$ and ensures the computation is tractable:
$\mathcal{L}_{\text{DSE}}$ $=\mathop{\mathbb{E}}_{\begin{subarray}{c}x_{0}\sim p_{0},\\ x\sim p(\cdot\mid x_{0})\end{subarray}}\bigg{[}\sum_{z\neq x}w_{xz}\Big{(}s_{\theta}(x)_{z}$
$\quad-\frac{p(z\mid x_{0})}{p(x\mid x_{0})}\log s_{\theta}(x)_{z}\Big{)}\bigg{]}$ (6)
where $p$ is assumed to be perturbation of some base density $p_{0}$ and weights $w_{xz}>0$.
The ELBO (Theorem 3.6 in Lou et al. (2024)) provides an upper bound on the negative log-likelihood, which is what we optimize for in generative models:
$-\log p^{\theta}_{0}(x_{0})\leq\mathcal{L}_{DWDSE}(x_{0})+constant$ (7)
where $\mathcal{L}_{DWDSE}$ integrates $\mathcal{L}_{DSE}$ weighted by the forward diffusion matrix. At inference time, we mask the label token in the template $X_{i}$ and use the model to predict it, restricting the possible outputs to valid labels in $\mathcal{V}_{\mathcal{Y}}$. For further details, refer to Lou et al. (2024).
4 Experiments
Our experiments are designed towards addressing the following research questions:
  1. Q1.
    How do different modeling approaches compare against each other when trained from scratch?
  2. Q2.
    How much does noise perturbation via random token substitution and token dropping affect the performance of different modeling approaches ?
  3. Q3.
    How well are the different modeling approaches calibrated ? For ordinal classification, how well the predicted distributions over ordinal categories follow a unimodal shape ?
4.1 Datasets
We evaluate our models on 9 text classification benchmark datasets to ensure a comprehensive assessment across multiple domains, text lengths, and classification types - sentiment analysis, movie reviews, news categorization, and social media analysis. These are: AG News Zhang et al. (2015), Emotion Saravia et al. (2018), Stanford Sentiment Treebank (SST2 & SST5) Socher et al. (2013), Multiclass Sentiment Analysis, Twitter Financial News Sentiment, IMDb Maas et al. (2011), and Hate Speech Offensive Davidson et al. (2017). These datasets encompass varying levels of complexity, ranging from binary text classification to fine-grained multi-class categorization, with textual inputs spanning from concise single sentences to extensive paragraph-level passages. Further details are postponed to Appendix C.
4.2 Experimental Setup
We conduct an extensive benchmarking study comparing the five different modeling approaches for text classification summarised in Section 3: AR, ARpseudo , MLM, DIFF, and ENC. These models are evaluated on 9 popular classification benchmark datasets as mentioned in Section 4.1.
[…]
In the second part of our evaluation, we assess each model’s robustness to input perturbations. In real-world scenarios, particularly in e-commerce platforms, often encounter various text corruptions (like OCR errors in product documentation, truncated reviews, or incomplete user queries), we focus on two systematic types of synthetic noise to evaluate model robustness: (a) Random Token Drop — where Xof tokens are randomly removed from the input sentence, and (b) Random Token Substitution — where Xof tokens are replaced with random tokens from the vocabulary (excluding special tokens like [PAD], [MASK]). We conduct these experiments to explore Q2 about model robustness to input perturbations.
Lastly, we assess model performance on calibration and ordinal metrics—aspects often overlooked but critical for real-world deployment. While a similar study exists for pre-trained models (Kasa et al., 2024), ours focuses on models trained from scratch. For ordinal classification tasks, we verify ordinal alignment, ensuring that the predicted probability distribution reflects the natural ordering of categories (e.g., the probability of “good” should be closer to “neutral” than to “bad”).
For ordinal evaluation, we report MSE (Mean Squared Error), MAE (Mean Absolute Error), and Unimodality (UM). For calibration, we measure ECE (Expected Calibration Error) and MCE (Maximum Calibration Error). UM verifies that the predicted probability distribution has a single peak, thus preserving class ordering—for instance, preventing models from assigning high confidence to both extremely positive and negative sentiments simultaneously. Calibration metrics quantify the discrepancy between predicted probabilities and empirical frequencies. For detailed descriptions, see Kasa et al. (2024) and Wang (2023). These experiments address Q3. Refer to Appendix B for details on hyperparameters and training setup.
Figure 2: [Best viewed in color] Comparison of weighted-F1 scores of models across different configurations ($\uparrow$ is better). For rest of the datasets, refer to Figure 8 in Appendix E. (X-axis: sample size, Y-axis: weighted-F1 score)
Figure 3: [Best viewed in color] Comparison of weighted-F1 scores between ARpseudo and AR ($\uparrow$ is better). 1-layer results are omitted here as they are mostly trivial in low-data settings. Results for remaining datasets are provided in Figure 9, Appendix E. (X-axis: sample size, Y-axis: weighted-F1 score)
5 Results
Figure 4: [Best viewed in color] Calibration and Ordinal performance of 12-layers model on SST-5. For ECE, MCE, MAE, MSE ($\downarrow$ is better) and UM ($\uparrow$ is better) (X-axis: sample size).
Config ENC ARpseudo AR MLM DIFF
(12L,12H) 51.1% 46.7% 37.8% 34.4% 54.4%
(6L,6H) 47.8% 51.1% 46.7% 46.7% 53.3%
Table 1: Minimum noisefor 10% $\Delta$ weighted-F1 drop under Random Token Dropping. ($\uparrow$ is better)
Config ENC ARpseudo AR MLM DIFF
(12L,12H) 32.2% 22.2% 28.9% 31.1% 35.6%
(6L,6H) 37.8% 21.1% 30.0% 32.2% 34.4%
Table 2: Minimum noisefor 10% $\Delta$ weighted-F1 drop under Random Token Substitution. ($\uparrow$ is better)
We analyze the results from all the experiments and provide valuable insights & recommendations for model selection.
Q1: For 1-layer, 1-head models (Figure 2), all approaches show near-random performance in low-data regimes. However, as training data increases, only ENC (orange line) continues to improve, ultimately outperforming others in high-data settings. This suggests that for small models - often necessary due to real-world latency constraints - ENC is the most effective approach. The classical ‘two regimes’ phenomenon does not manifest when the model size is small.
The pattern shifts dramatically for larger architectures.Under the 12-layer, 12-head configuration, both generative models—AR and DIFF—outperform ENC in low-data settings, with this advantage diminishing as data increases. This aligns with previous findings (Ng and Jordan, 2001; Yogatama et al., 2017; Rezaee et al., 2021) about generative models’ advantages in data-limited scenarios. Surprisingly, for large models, the pseudo-generative MLM (blue line) consistently outperforms all methods across our 9 benchmark datasets in high-data settings, challenging the conventional wisdom about discriminative dominance in high-sample regime. This aligns with Erhan et al. (2010)’s finding that pseudo-generative models implicitly perform unsupervised pre-training alongside supervised learning, creating an effective multi-task setup (Section 2). Their work shows that this unsupervised phase acts as a data-dependent regularizer, guiding optimization toward better-generalizing minima. For large models, direct fine-tuning without this implicit pre-training often leads to suboptimal convergence, explaining ENC’s underperformance relative to MLM. Thus, for scenarios without model size constraints, generative models emerge as the optimal choice for low-data settings such as for low-resource languages and continual learning applications requiring frequent updates with limited samples, while pseudo-generative MLM is superior when abundant labeled data is available.
Another noteworthy observation is that under the 6-layer, 6-head configuration, in low-data settings, DIFF emerges as the best performing model across all datasets, clearly outperform even it’s generative counterpart AR. As the training data size increases, we see that the discriminative ENC outperforming DIFF. Thus, in medium scale architectures, between the generative DIFF and the discriminative ENC, the classical ‘two regimes’ still holds.
Figure 3 shows that ARpseudo generally underperforms AR and also displays higher variance in low-data settings—the recommended use case—while the opposite holds in high-data scenarios. This reveals a new insight beyond Li et al. (2025), who only evaluated full-data settings where ARpseudo performed better in-distribution. As noted in Section 3, AR requires $|label|$-times forward passes per prediction, unlike the single pass needed for ARpseudo; however, this can be mitigated via batching or parallel processing, reducing inference time differences at the cost of higher computation.
Q2: We evaluate the robustness of all approaches under both 6-layer and 12-layer configurations across two noise schedules in full-data settings. We exclude 1-layer models from this analysis since their performance is mostly trivial (except for ENC), making robustness comparisons uninformative. In Tables 1 and 2, we report the minimum noise level at which each model’s performance drops by 10% $\Delta$ (similar trend for 5% & 15%) relative to its peak, averaged across all datasets, as a measure of robustness boundary. Our analysis reveals that all models exhibit lower robustness to substitution noise compared to dropping. This can be explained by the inequality: $P(\text{garbage}_{t}|X_{1...t-1})<P(x_{t+1}|X_{1...t-1})$—the model is more likely to assign lower probability to a corrupted token than to a skip token at $t+1$-th position (assuming $t$-th token was dropped), which may still be contextually relevant given $X_{1...t-1}$.
The generative DIFF demonstrates superior robustness to both token dropping and substitution (except in 6-layers where ENC is slightly better), likely because its training paradigm involves recovering true tokens from noise/masked inputs. The discriminative ENC maintains consistent robustness under both noise types, while generative AR shows the high sensitivity to noise. Combining these findings with Q1’s results reveals that generative AR models face dual challenges compared to ENC in full data settings: they underperform in terms of both weighted-f1 and robustness. This contrasts with Li et al. (2025)’s findings that discriminative ENC models rely on shortcuts and show less robustness compared to generative AR. However, their analysis focuses on shortcut learning and distribution shifts rather than input perturbation noise across varying model sizes. Notably, while the pseudo-generative MLM and ARpseudo demonstrates superior performance in larger models at full data settings, they exhibit lower robustness compared to similarly performing ENC models. Moreover the relative drop in robustness in moving from dropping to substitution noise is more severe in ARpseudo compared to AR. This is likely because ARpseudo conditions on corrupted inputs, so it’s directly affected by garbage tokens polluting the predictive context. However, for AR clean label conditions the model, and the noisy input is scored globally — giving the model more flexibility to discount garbage.
Q3: Figure 4 presents ordinal and calibration results for SST-5, selected for its balanced distribution, inherent class ranking (e.g., very positive to very negative), and highest number of classes. Results for other datasets are in Appendix D. DIFF does not support calibration metrics like ECE, MCE, and UM, as its masking/absorbing noise process produces only binary outputs rather than soft probabilities. While a uniform noise schedule can yield probabilities over $\mathcal{V}$, it performed slightly worse, so we used the absorbing schedule in our study.
From the ECE and MCE plots, we observe that ENC outputs remain well-calibrated across all sample sizes, while MLM reaches similar calibration only in high-data regimes. We also see that MLM and ENC achieve UM in over 80% of the samples, aligning with findings from Kasa et al. (2024). Their MAE and MSE values are also low, indicating strong ordinality in high-data settings. This completes the picture for large models under high-data, where MLM not only outperforms others in weighted-F1 but is also well-calibrated and ordinal, making it a strong candidate for real-world deployment. However, under low-data conditions, 12-layers AR outperforms ARpseudo in 7 out of 9 datasets on calibration metrics. It also surpasses DIFF in ordinal performance, thus making it the more reliable choice among generative models in low-data scenario. Also, even though generative approaches like DIFF were recommended earlier in Q1 based on weighted-F1 for 6-layers case (in Figure 2) deploying them in production could be risky when calibrated or ordinal probabilities are required, especially for imbalanced datasets like twitter and hatespeech (see Appendix D). These metrics are particularly important when downstream models consume output probability scores as features which is often the case in multi-stage ranking systems.
Lastly, Figure 5 in Appendix D reveals an interesting trend: as model size increases, calibration metrics either remain flat or worsen. This suggests that larger models or improved classification accuracy do not necessarily lead to better calibration, aligning with the findings of Guo et al. (2017) where they show similar behaviour using ResNets (He et al., 2016). However, for ordinal metrics, we observe substantial improvements when moving from 1-layer to 6-layer models, with performance plateauing at 12 layers. A similar trend was reported in Kasa et al. (2024) for pre-trained models.
6 Conclusion
Our study offers practical modeling recommendations across deployment scenarios. For latency-sensitive applications, ENC is ideal—especially in the 1-layer setting—due to its efficiency, robustness to noise, and well-calibrated, ordinal outputs. For offline settings with sufficient data, the 12-layer MLM performs best across F1, calibration, and ordinal metrics, though caution is needed with noisy inputs due to its lower robustness to token dropping. In low-resource scenarios, both AR and DIFF are strong options, with DIFF favored for its noise resilience and performance at 6-layers. However, if calibrated probability outputs are essential, such as in ranking pipelines, AR is the preferred choice.
7 Limitations
While we conducted a thorough examination of generative and discriminative classifiers under standard i.i.d. assumptions, our findings may not generalize to scenarios involving distribution shifts, such as co-variate shift (Bickel et al., 2009) or concept shift (Roychowdhury et al., 2024). Our analysis was limited to traditional fine-tuning approaches, excluding emerging paradigms such as few-shot prompt-based in-context learning (Sun et al., 2023; Gupta et al., 2023) and parameter-efficient techniques like LoRA (Hu et al., 2022), which may uncover newer insights. Furthermore, our study focused exclusively on pure text classification, leaving the exploration of multi-modal scenarios involving tabular data (Pattisapu et al., 2025), images Lu et al. (2019), audio Kushwaha and Fuentes (2023), and other modalities for future work. An additional promising direction would be to investigate the use of pre-trained models, disentangling the effects of pre-training and fine-tuning (see Section 1), and to assess their effectiveness in cross-lingual transfer tasks. We leave these avenues for future exploration.
Appendix A More Background and Related works
Discrete Diffusion Models for Classification. Recent advances in discrete diffusion models have shown promising results in text generation tasks, matching or surpassing autoregressive models at GPT-2 scale (Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024). While these models have demonstrated success in controlled generation tasks (Li et al., 2022a; He et al., 2023), specifically syntax controlled generation of text Kumar et al. (2020) and text infilling, their application to classification remains relatively unexplored. Traditional diffusion models for text generation, such as DiffusionBERT (He et al., 2023), DiffusionLM (Li et al., 2022b), and D3PM (Austin et al., 2021), operate by embedding discrete token sequences into continuous spaces and applying Gaussian noise-based diffusion. In contrast, SEDD (Lou et al., 2024) was the first to directly model diffusion in discrete space through a score entropy-driven objective. Hence, we adopt SEDD as our baseline method. Our work provides the first systematic evaluation of discrete diffusion models for classification tasks, comparing them against traditional discriminative and generative approaches.
Robustness to Noise. Previous studies have examined robustness primarily through the lens of adversarial attacks Li et al. (2019), distribution shifts Li et al. (2025) and domain shifts Jaini et al. (2024). While recent work has provided certified robustness guarantees for perturbations like insertion, deletion, reordering and synonyms for specific architectures Zeng et al. (2023); Zhang et al. (2024), our study presents comparisons across model families under two different noise conditions in the context of TC for transformer architectures.
Calibration & Ordinality. Model calibration is crucial in classification, as it reflects how well predicted probabilities align with actual frequencies. Proper Scoring Rules (PSR) Merkle and Steyvers (2013) offer a theoretical basis for producing calibrated predictions: a scoring rule (i.e. loss function) is proper if its expected value is minimized only when predicted probabilities match the true distribution. All our modeling approaches—Generative (AR, MLM, Discrete Diffusion) and Discriminative (Encoder)—optimize proper scoring rules. GPT and MLM maximize likelihood, Discrete Diffusion optimizes a variational bound, and cross-entropy minimizes the KL-divergence between predicted and true distributions. Recent work Blasiok et al. (2023) shows that models trained with PSRs are often naturally calibrated when achieving low training loss, without requiring post-hoc calibration. This motivates us to empirically assess calibration across our models, as their differing architectures and objectives may still lead to varying calibration behaviors.
Ordinality in text classification is essential for applications like sentiment analysis or medical assessments, where label order affects decisions and distant misclassifications are more harmful. Recent works Kasa et al. (2024) systematically compare explicit methods—like custom losses enforcing label order—with implicit approaches using pretrained models’ semantics. However, no prior work focuses on exploring ordinality across diverse modeling frameworks trained from scratch.
Appendix B Implementation Details
We use the bert-base-uncased [^1] https://huggingface.co/google-bert/bert-base-uncased architecture as the backbone for our Encoder and MLM experiments, without initializing the model with pretrained weights. This architecture contains approximately 110M parameters, comprising 12 encoder layers, 12 attention heads, and a hidden size of 768. We run all experiments for 3 random seeds and report the average and standard deviation results in main paper.
For the Encoder experiments, we conducted a grid search over several hyperparameters, including learning rates of {1e-5, 2e-5, 3e-5, 4e-5, 5e-5}, batch sizes of {32, 64, 128, 256}, and a fixed sequence length of 512 tokens. Training was performed for 30 epochs uniformly for all datasets without early stopping. For the MLM-based experiments, we retained similar hyperparameter ranges but trained for 200 epochs to account for the increased complexity of masked token prediction. We observed that adding an early stopping patience parameter sometimes led the model to select a suboptimal checkpoint, as the validation loss often continued to decrease gradually after remaining flat or oscillating for several epochs.
For the AR and ARpseudo experiments, we used the GPT-2 base architecture [^2] https://huggingface.co/openai-community/gpt2 as the backbone with 137M parameters comparable with our other experiments. We trained a causal language model to minimize the next-token prediction loss over the concatenated input and label sequence. A grid search was conducted with the same hyperparameter range as mentioned above. The models were trained for up to 100 epochs, with early stopping based on validation loss, using a patience parameter of 10 epochs.
Our Text Diffusion approach follows the Diffusion Transformer architecture Peebles and Xie (2023) which is basically the vanilla transformer encoder with an extra time-conditioned embedding incorporated with it. The parameter count is $\sim$160M due to the addition of time-dependent embeddings required by the diffusion mechanism. To counter this, we conducted an ablation study by increasing the encoder size to 160M parameters (by adding layers) for other approaches (like ENC, MLM) to match the diffusion model size, but observed no difference in performance. Hence we retain their original settings as reported above. For diffusion-specific hyperparameters, we used a batch size of 64, learning rate 3e-4 and trained for 200K iterations. We adopted a geometric noise schedule that interpolates between $10^{-4}$ and 20, similar to the setup in Lou et al. (2024), and used the following absorbing/masking matrix $Q^{absorb}$ as part of the transition modeling. This was the best hyperparameter setting we found.
$Q_{\text{absorb}}=\begin{bmatrix}-1&0&\cdots&0\\ 0&-1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&0\end{bmatrix}$
All experiments were conducted using multi-GPU training across eight NVIDIA A100 GPUs. Training time varied depending on the methods and configurations used for each dataset. The range of training times (in hours) for various datasets is presented in Table 3. All reported training times correspond to full-data training configurations.
Config ENC ARpseudo AR MLM DIFF
(1L,1H) 1-2 2-4 2-4 1-4 1-4
(6L,6H) 1-3 3-7 3-7 3-7 2-6
(12L,12H) 2-5 5-10 5-10 5-10 5-12
Table 3: Training time (in hrs) ranges across different datasets for each configuration and approach.
Appendix C Dataset Details
Dataset Split Examples Classes Avg Tokens Label Dist. (%) Ordinal
IMDb train 25,000 2 313.87 0: 50.0, 1: 50.0 $\times$
test 25,000 2 306.77 0: 50.0, 1: 50.0
agnews train 120,000 4 53.17 0-3: 25.0 each $\times$
test 7,600 4 52.75 0-3: 25.0 each
emotion train 16,000 6 22.26 0: 29.2, 1: 33.5, 2: 8.2, 3: 13.5, 4: 12.1, 5: 3.6 $\times$
test 2,000 6 21.90 0: 27.5, 1: 35.2, 2: 8.9, 3: 13.8, 4: 10.6, 5: 4.1
hatespeech train 22,783 3 30.04 0: 5.8, 1: 77.5, 2: 16.7
test 2,000 3 30.18 0: 5.5, 1: 76.6, 2: 17.9
multiclasssentiment train 31,232 3 26.59 0: 29.2, 1: 37.3, 2: 33.6
test 5,205 3 26.91 0: 29.2, 1: 37.0, 2: 33.8
rottentomatoes train 8,530 2 27.37 0: 50.0, 1: 50.0 $\times$
test 1,066 2 27.32 0: 50.0, 1: 50.0
sst2 train 6,920 2 25.21 0: 47.8, 1: 52.2 $\times$
test 872 2 25.47 0: 49.1, 1: 50.9
sst5 train 8,544 5 25.04 0: 12.8, 1: 26.0, 2: 19.0, 3: 27.2, 4: 15.1
test 1,101 5 25.24 0: 12.6, 1: 26.3, 2: 20.8, 3: 25.3, 4: 15.0
twitter train 9,543 3 27.62 0: 15.1, 1: 20.2, 2: 64.7
test 2,388 3 27.92 0: 14.5, 1: 19.9, 2: 65.6
Table 4: Dataset statistics showing training and test split sizes, number of classes, mean and maximum token lengths, and label distribution percentages. Refer to Section C for details on datasets.
AG News Zhang et al. (2015): It consists of approximately 120K training samples and 7.6K test samples, divided into four categories: World, Sports, Business, and Technology. Each sample contains a short news article, typically consisting of the title and the first few sentences. Emotion Saravia et al. (2018): A collection of English tweets labeled with six basic emotions: anger, fear, joy, love, sadness, and surprise. It is designed for emotion detection in text. The dataset has 20K samples divided into 16K samples for training and 2K samples each for validation and testing. Stanford Sentiment Treebank (SST) Socher et al. (2013): We utilize both the SST-2 (binary sentiment) and SST-5 (fine-grained sentiment) variants of the Stanford Sentiment Treebank dataset. SST-2 consists of sentences labeled as either positive or negative, suitable for binary sentiment classification, while SST-5 includes five sentiment categories: very negative, negative, neutral, positive, and very positive, allowing for more fine-grained sentiment analysis. Multiclass Sentiment Analysis  [^3] https://huggingface.co/datasets/Sp1786/multiclass-sentiment-analysis-dataset: This dataset consists of 41.6K data points, labeled into three sentiment categories: positive, negative, and neutral. While the dataset is designed for multiclass sentiment classification, it exhibits class imbalance, with certain sentiment classes being more prevalent than others. This imbalance provides a more realistic challenge for sentiment analysis models, testing their ability to handle skewed distributions and still perform effectively across all sentiment categories. Twitter Financial News Sentiment  [^4] https://huggingface.co/datasets/zeroshot/twitter-financial-news-sentiment: A specialized English-language collection of finance-related tweets, annotated for sentiment analysis. It consists of 11,932 tweets labeled with three sentiment categories: Bearish, Bullish, and Neutral. This dataset is designed to test models’ ability to understand domain-specific language and nuanced sentiment expressions in financial contexts. IMDb Maas et al. (2011): A binary sentiment analysis dataset consisting of 50K reviews from the Internet Movie Database (IMDb), labeled as positive or negative. The dataset is balanced, with an equal number of positive and negative reviews. This dataset is characterized by longer document lengths and detailed opinions, making it a challenging benchmark. Rotten Tomatoes Pang and Lee (2005): A binary classification dataset which contains 10,662 movie review sentences, equally divided into 5,331 positive and 5,331 negative examples. The dataset is characterized by relatively short, opinion-driven sentences that reflect concise sentiments about films. Hate Speech Offensive Davidson et al. (2017): A major challenge in automatic hate speech detection is distinguishing hate speech from other forms of offensive language. This dataset consists of approximately 25K tweets, labeled into three categories: hate speech, offensive language without hate speech, and neutral content.
Refer to Table 4 for details on dataset statistics.
Appendix D More Ordinal & Calibration Results
In this section, we take a closer look at ordinal and calibration results for the datasets decribed above. Here we report ordinal metrics on the datasets Stanford Sentiment Treebank (SST5) Socher et al. (2013), Multiclass Sentiment Analysis, Hate Speech Offensive Davidson et al. (2017) and Twitter Financial News Sentiment since these are the only multi-class ordinal datasets out of 9. Calibration metrics are reported on all 9 datasets.
In Figure 5, we compare how ordinal and calibration metrics vary with increasing model size. Figure 6 presents the ordinal metrics for all four ordinal datasets, while Figure 7 shows the calibration metrics for all nine datasets. The corresponding insights are discussed in Section 5 (see Q3).
Figure 5: [Best viewed in color] Calibration and Ordinal metrics comparison across layers 1, 6 and 12. For ECE, MCE, MAE, MSE, ($\downarrow$ is better) and UM ($\uparrow$ is better).
Figure 6: [Best viewed in color] Ordinal metrics. For MAE, MSE, ($\downarrow$ is better) and UM ($\uparrow$ is better).
Figure 7: [Best viewed in color] Calibration metrics. For ECE, MCE ($\downarrow$ is better)
Appendix E More Main Results
This section contains the extended results of Figure 2 (see Figure 8) and Figure 3 (see Figure 9) for all 9 datasets. We omit 1-layer plots for Figure 9 since the performance is mostly trivial for low-data settings and the same trend is observed as 6/12-layers for full-data settings.
Figure 8: [Best viewed in color] Comparison of weighted-F1 scores of models across different configurations for all 9 datasets. ($\uparrow$ is better) (X-axis: sample size, Y-axis: weighted-F1 score)
Figure 9: [Best viewed in color] Comparison of weighted-F1 scores between ARpseudo and AR ($\uparrow$ is better) for all datasets. (X-axis: sample size, Y-axis: weighted-F1 score)